OpenStudy (nhra34x0):

A hose fills a tank at a rate of 30 gallons per minute. What is the rate in liters per hour? A hippo can travel up to 40 miles per hour. What is the hippo’s speed, to the nearest tenth, in feet per minute?

OpenStudy (sapphiremoon):

A couple of conversions we can do with the factor label (or conversion factor) method. First, to get 30gal/min to L/h, we can convert minutes to hours, then gallons to liters. \[\frac{30 gal}{\min} \times \frac{60 \min}{h} \times \frac{3.79 L}{gal}\] Gallons and minutes cancel, leaving us with 30 x 60 x 3.79 = 6822 L/h. Big, but if we have a lot of gallons per minute, we should expect more liters per hour. To get 40 mi/h to ft/min, it's much the same thing, miles to feet and hours to minutes: \[\frac{40 mi}{h} \times \frac{5280 ft}{mi} \times \frac{1 h}{60 \min}\] Miles and hours cancel to give us 40 x 5280 /60 = 3520 ft/min. Also within reason. Make sense?

OpenStudy (nhra34x0):

Yes, thank you .